# Masked Particle Modeling on Sets: Towards Self-Supervised High Energy Physics Foundation Models

### v1

@@ V.3 Fine-Tuning for Jet Classification
Figure 3: Accuracy of different training strategies as a function of the number of labelled training samples. Accuracy is calculated on the ten classes in the JetClass dataset. The average and standard deviation over 5 trainings is shown in solid lines and uncertainty bands, respectively.
@@ V.4 Fine-Tuning for Jet Classification on New Classes
To test whether the pre-trained model learns features that are generically useful for downstream tasks, or only those which are useful for the classes that are seen in the pre-training set, we perform an out-of-context test where we pre-train on a subset of the classes provided in JetClass and test the fine-tuning on the remaining classes. Specifically, the pre-training is performed on the six Higgs and QCD classes, and then fine-tuned on the remaining four classes. The results shown in Fig. 4 indicate that both pre-trained models, with fixed and with fine-tuned backbone, outperform the fully supervised training when only a limited amount of data is available. This indicates that even with small amounts of labelled data, the representations learned during pre-training are generically useful for out-of-context classes. As also expected, with enough labelled training data, the fully supervised model can surpass the fixed backbone model and converge to a similar performance to the fine-tuned model. In essence, the additional labelled data allows both the fine-tuned backbone model and the supervised model to adapt the representations to the specific downstream tasks.
Figure 4: Accuracy of different training strategies as a function of labelled training samples. Accuracy is calculated on four classes, held out from all pre-trained models, in the JetClass dataset. The average and standard deviation over 5 trainings is shown in solid lines and uncertainty bands, respectively.
@@ V.5 Fine-Tuning for Jet Classification on New Data Sets
We also test how well the backbone representation works out-of-domain on a different dataset of jets and how well the model can adapt to this new domain with fine-tuning. In this case, pre-training is performed with one dataset, JetClass, and fine-tuning is performed with a different data set, RODEM. More abstractly, this is a test of how well such a pre-training and fine-tuning strategy with mixed data sets may help address the domain shift challenge in HEP. This challenge come from the small, but potentially significant, differences between simulated data and real experimental data, such that models trained on simulated data have differences in prediction performance between simulated and real data, thus causing systematic uncertainties. With the MPM scheme, we can pre-train directly on unlabelled data, i.e. we can pre-train directly on real experimental data for representation learning. Thus we can also explore how well the features learned from data may maintain predictive power even after fine-tuning on simulation, represented here by the second labelled data set used for fine-tuning. In general, training primarily on real data has many potential benefits. It promises a reduced sensitivity to simulation-related domain shifts and calibration effects, leading to an overall simplified calibration procedure. An overall reduced simulation budget may also help alleviate the growing compute and storage limitations of the LHC experiments.
@@
Figure 5: The QCD rejection at $50\%$ top-jet efficiency evaluated on (left) the RODEM test set and (right) the JetClass test set, as a function of the size of the RODEM data set used for fine-tuning. All models are pre-trained on JetClass. The average and standard deviation of rejection over 5 trainings is shown in solid lines and uncertainty bands, respectively.
@@ V.6 Weakly supervised classification
Figure 6: Models trained with weak supervision to classify data sets of different label proportions. A data set of one million QCD jets is compared to a data set with one million QCD jets plus $N$ top jets. The average and standard deviation of the significant improvement over 5 trainings is shown in solid lines and uncertainty bands, respectively.

### v3

@@ V.3 Fine-Tuning for Jet Classification
Figure 3: Accuracy of different training strategies as a function of the number of labelled training samples. Accuracy is calculated on the ten classes in the JetClass dataset. The average and standard deviation over 5 trainings is shown in solid lines and uncertainty bands, respectively. Models with frozen pretrained backbone weights during fine-tuning are “Fixed”, and those with updated weights are “Fine-tuned”.
@@ V.4 Fine-Tuning for Jet Classification on New Classes
To test whether the pre-trained model learns features that are generically useful for downstream tasks, or only those which are useful for the classes that are seen in the pre-training set, we perform an out-of-context test where we pre-train on a subset of the classes provided in JetClass and test the fine-tuning on the remaining classes. Specifically, the pre-training is performed on the six Higgs and QCD classes, and then fine-tuned on the remaining four classes. The four remaining classes are top quarks decaying with and without a lepton, and $W,Z$ decays each as a distinct class. The results shown in Fig. 4 indicate that both pre-trained models, with fixed and with fine-tuned backbone, outperform the fully supervised training when only a limited amount of data is available. This indicates that even with small amounts of labelled data, the representations learned during pre-training are generically useful for out-of-context classes. As also expected, with enough labelled training data, the fully supervised model can surpass the fixed backbone model and converge to a similar performance to the fine-tuned model. In essence, the additional labelled data allows both the fine-tuned backbone model and the supervised model to adapt the representations to the specific downstream tasks.
Figure 4: Accuracy of different training strategies as a function of labelled training samples. Accuracy is calculated on four classes, held out from all pre-trained models, in the JetClass dataset. The average and standard deviation over 5 trainings is shown in solid lines and uncertainty bands, respectively. Models with frozen pretrained backbone weights during fine-tuning are “Fixed”, and those with updated weights are “Fine-tuned”.
@@ V.5 Fine-Tuning for Jet Classification on New Data Sets
We also test how well the backbone representation works out-of-domain on a different dataset of jets and how well the model can adapt to this new domain with fine-tuning. In this case, pre-training is performed with one dataset, JetClass, and fine-tuning is performed with a different data set, RODEM. More abstractly, this is a test of how well such a pre-training and fine-tuning strategy with mixed data sets may help address the domain shift challenge in HEP. This challenge come from the small, but potentially significant, differences between simulated data and real experimental data, such that models trained on simulated data have differences in prediction performance between simulated and real data, thus causing systematic uncertainties. As such, this test is in analogy with the idea that one may want to pre-train on real data, and only fine-tune on small simulated data sets, and we hope to explore the applicability of such an approach. With the MPM scheme, we can pre-train directly on unlabelled data, i.e. we can pre-train directly on real experimental data for representation learning. Thus we can also explore how well the features learned from data may maintain predictive power even after fine-tuning on simulation, represented here by the second labelled data set used for fine-tuning. In general, training primarily on real data has many potential benefits. It promises a reduced sensitivity to simulation-related domain shifts and calibration effects, leading to an overall simplified calibration procedure. An overall reduced simulation budget may also help alleviate the growing compute and storage limitations of the LHC experiments.
@@
Figure 5: The QCD rejection at $50\%$ top-jet efficiency evaluated on (left) the RODEM test set and (right) the JetClass test set, as a function of the size of the RODEM data set used for fine-tuning. All models are pre-trained on JetClass. The average and standard deviation of rejection over 5 trainings is shown in solid lines and uncertainty bands, respectively. Models with frozen pretrained backbone weights during fine-tuning are “Fixed”, and those with updated weights are “Fine-tuned”.
@@ V.6 Weakly supervised classification
Figure 6: Models trained with weak supervision to classify data sets of different label proportions. A data set of one million QCD jets is compared to a data set with one million QCD jets plus $N$ top jets. The average and standard deviation of the significant improvement over 5 trainings is shown in solid lines and uncertainty bands, respectively. Models with frozen pretrained backbone weights during fine-tuning are “Fixed”, and those with updated weights are “Fine-tuned”.